OpenStudy (anonymous):

Explain the purpose of converting a radical expression into a rational exponent.

OpenStudy (anonymous):

It can make the arithmetic between radical expressions easier

OpenStudy (anonymous):

can you give me an easy example or two?

OpenStudy (anonymous):

\[ \Large \sqrt{3^4} = 3^{4/2} = 3^2 = 9 \]

OpenStudy (anonymous):

If you have: \[\sqrt[3]{23}^{4}*\sqrt[4]{23}^{5}\]That is the same as:\[23^{\frac{ 4 }{ 3 }}*23^{\frac{ 4 }{ 5 }}=23^{\frac{ 32 }{ 15 }}\]

OpenStudy (anonymous):

The first one is very hard to solve in that form but once you put it into rational exponent form, you can multiply them easily. @wio example was also good